OpenStudy (anonymous):

If y=9 when x=3, find x when y=7

OpenStudy (anonymous):

Clearly, the answer must be \[ \frac{\sqrt{\pi}}{\gamma^2} \]

OpenStudy (anonymous):

No, seriously you need more data to predict values of x ;).

OpenStudy (anonymous):

x=1